OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−4, 8) and (−3, −6). 7 −14 0 14

OpenStudy (callisto):

similar to previous one, use the formula and try to do it once before you ask

OpenStudy (anonymous):

−14 isn't?

OpenStudy (callisto):

slope = (y2 -y1)/(x2- x1)

OpenStudy (callisto):

slope = (y2 -y1)/(x2- x1)= [8-(-6)]/ [(-4)-(-3)] = 14/-1 =-14